OpenStudy (anonymous):

Easy math Match the letters with the matching number. 1)$42 over 2 years at an interest rate of 3% 2)$40 over 4 years at an interest rate of 2% 3)$48 over 2 years at an interest rate of 4% 4)$40 over 2 years at an interest rate of 5% A)p=$500 B)p=$600 C)p=$700

OpenStudy (amistre64):

good thing i aint getting a grade on it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

but i am

OpenStudy (amistre64):

what type of interest? simple, compounded, continuous?

OpenStudy (anonymous):

simple

OpenStudy (amistre64):

well then, percents are a measure of 100 parts; so we have to divide our number by 100. for example: 42/100 = 1% of 42 ; but we want 3 of them 42(3)/100 = 3%, simple interest for 1 year

OpenStudy (amistre64):

i dont see how the answers are the interest; unless there is some other part of the instructions that you are just not telling

OpenStudy (anonymous):

Match the information about interest earned with the principal amount invested. Remember the equation for calculating simple interest is .

OpenStudy (amistre64):

ooooo .... so your given interest and match it to prinipal

OpenStudy (anonymous):

I=prt

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

so what parts got you stumped?

OpenStudy (anonymous):

sorry i left that out its very important

OpenStudy (anonymous):

all of it

OpenStudy (amistre64):

well, the problem suggests that youve already been taught "most" of it at least

OpenStudy (amistre64):

and i dont really have the time to reinvent the wheel on it

OpenStudy (amistre64):

If we know the formula; we just fill in the parts to find whats missing

OpenStudy (amistre64):

I = p r t for $42(I) over 2(t) years at an interest rate of 3% (r) 42 = p .03 2 42 ----- = p 2*.03

OpenStudy (anonymous):

the formula is I=prt but i get confused

OpenStudy (amistre64):

rearrange it to fit better; solve it for "p" I/(rt) = p

OpenStudy (anonymous):

oh ok thanks so can you match the letters and numbers together please.

OpenStudy (amistre64):

i could yes, but id rather dbl chk your efforts; that way i dont feel like im just being used for cheap labor. You show me yours and all show you mine :)

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

1)$42 over 2 years at an interest rate of 3% 2)$40 over 4 years at an interest rate of 2% 3)$48 over 2 years at an interest rate of 4% 4)$40 over 2 years at an interest rate of 5% this is so i aint got to keep looking up top for it

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

Well, mine are ready :)

OpenStudy (anonymous):

so what do you want me to do now.

OpenStudy (amistre64):

show me what your best effort is; right or wrong, show me what you get. then ill post mine for you to compare to.

OpenStudy (anonymous):

ok i think 1) goes with C 2) A 3)B 4)D

OpenStudy (amistre64):

fair enough :) 1)$42 over 2 years at an interest rate of 3% 42/(2*.03) = 700 2)$40 over 4 years at an interest rate of 2% 40/(4*.02) = 500 3)$48 over 2 years at an interest rate of 4% 48/(2*.04) = 600 4)$40 over 2 years at an interest rate of 5% 40/(2*.05) = 400

OpenStudy (amistre64):

i believe we agree

OpenStudy (anonymous):

yea thanks for the help!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank you